# revtex4-2Repair the float

### v1

@@ Abstract
We report temperature measurements using a transmon qubit by detecting the population of the first three levels of it, after employing a sequence of $\pi$-pulses and performing projective dispersive readout. We measure the effective temperature of the qubit and characterize its relaxation and coherence times $\tau_{1,2}$ for three devices in the temperature range 203002030020–30020 – 300 mK. Signal-to-noise (SNR) ratio of the temperature measurement depends strongly on $\tau_{1}$, which drops at higher temperatures due to quasiparticle excitations, adversely affecting the measurements and setting an upper bound of the dynamic temperature range of the thermometer. The measurement relies on coherent dynamics of the qubit during the $\pi$-pulses. The effective qubit temperature follows closely that of the cryostat in the range $100-250$ mK. We present a numerical model of the qubit population distribution and compare it favorably with the experimental results.
@@ I Introduction
In recent years theoretical proposals for quantum thermometry have attracted a lot of attention due to prospects of monitoring temperature with minimal invasion, properties of quantum systems with high accuracy at extremely low temperatures and opportunities for studying open quantum systems and quantum thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, the number of conceptually different experimental realizations of quantum thermometers based on coherent quantum objects remains quite limited. For example, quantum thermometry was demonstrated using a nuclear magnetic resonance (NMR) techniques for polarization measurements of spin ensembles in solutions at room temperature [17] and using coherent quantum N00N-states [18]. Generation and transfer of the latter was recently realized on the basis of a superconducting qubit [19]. Also there are noteworthy experiments on thermometry based on nitrogen-vacancy (NV) colour centres in diamond [20, 21, 22, 23, 24]. Another possible platform for quantum thermometers is semiconductor quantum dots, which were used for spectroscopy-type measurements in sub-kelvin temperature range in Ref. [25], exhibiting good signal-to-noise ratio and fast operation but limited coherence.
Prospects of utilizing superconducting qubits as thermometers are of dual research interest: they can facilitate studies of open quantum systems and interaction between the qubits and their environment, which is vital in the context of building large multiqubit systems such as quantum computers and simulators. On the other hand qubit thermometers can be used for realization of quantum thermometry protocols. Quantum electrodynamics (QED) circuits based on coherent superconducting transmon qubits were primarily used to measure the residual qubit population and the corresponding residual temperature in a number of works, for example in [26, 27, 28, 29]. The residual population of a transmon in a 3D cavity was measured by Rabi oscillations on the $e-f$ transition in Ref. [27] and correlation measurements in Ref. [28]. In Ref. [29] another technique of applying sequences of $\pi$-pulses was used to measure the effective qubit temperature by observing the population distribution for the three lowest energy states of a transmon in a planar geometry. A different approach was demonstrated in Ref. [30], where thermometry was done by reflection-type spectroscopy measurements of a transmon qubit strongly coupled to a waveguide.
@@
Yet qubits are sensitive to various sources of dissipation and decoherence, and the diverse nature of their environment (see, for example, [31, 32, 33, 34, 35, 36, 37, 38, 39]). Therefore in practice, identification of a specific thermal bath, leading to thermalization of the qubit, can be quite challenging; i.e., it may be difficult to identify what temperature the qubit is actually measuring.
@@
In this paper we present a comprehensive and concrete study of employing a superconducting qubit for quantum thermometry at sub-kelvin temperatures. The measurement technique used in our work relies on applying sequences of $\pi$-pulses for manipulating the qubit population distribution, requiring sufficient coherence of the qubit, and then performing consecutive qubit state readout, which was described in Ref. [29]. We also study temperature dependencies of qubit relaxation and coherence times. They determine the working range of the thermometer in terms of temperature, and they give conatraints on the pulse durations in the measurement. While showing a solid thermalization of the qubit to the cryostat stage, the technique allows us to measure the population distribution at temperatures up to $200-220$ mK, when the second excited state of the qubit was already significantly populated. This exceeds the applicability range of methods based on consideration of only the lowest qubit levels. At the same time the technique is less dependent on the signal-to-noise ratio (SNR) of the measurement setup. We compare the experimental results with a numerical model of qubit population measurements, being in a good qualitative agreement. Within the model we investigate the influence of the measured parameters on the population and effective qubit temperature obtained as the result. Finally, we address the factors limiting the dynamical range of a superconducting qubit thermometer. The paper is organized as follows: in Section II we provide a brief description of a realistic model of a qubit being in thermal equilibrium with heat baths. Section III explains the principle of the population distribution measurements. Section IV provides a brief description of the samples used in the experiment, and discussion on the main experimental results is given in Section V, and finally, limits of the thermometer and ways of mitigation are discussed in Section VI. The paper is followed by appendices providing additional experimental data, details of the numerical simulations, experimental setup and sample fabrication.
@@ II Qubit population and effective temperature
Let us consider a qubit in thermal equilibrium with a heat bath having temperature $T$. In this case the qubit is in a thermal state, which means that its density matrix is diagonal and the probabilities $p_{g}$ and $p_{e}$, respectively, to be in the ground |$\ket{g}$ or excited state |$\ket{e}$ (qubit population) are Boltzmann distributed as
@@
In a more realistic scenario when a qubit is in steady state with respect to several uncorrelated heat baths (with temperatures $T^{(i)}$, where $i$ enumerates baths), the state of the qubit is still a thermal-like state. But the temperature characterizing this state is not necessarily the temperature $T^{(i)}$ of any of the thermal baths. Then Eq. (1) could be interpreted as giving an effective qubit temperature $T_{\mathrm{eff}}$. In this sense, Eq. (1) can be rewritten as a definition of $T_{\mathrm{eff}}$ as
@@
which coincides with the true temperature in equilibrium.
For the case of ohmic heat baths, one can find the energy relaxation rate $\gamma_{1}$ for the qubit towards $T_{\mathrm{eff}}$ using Fermi golden rule approach. The detailed description is given in Appendix A. The qubit energy relaxation rate writes
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma_{1}^{(i)}\left[2n\left(T^{(i)},\omega_{ge}\right)+1\right]\\ =\sum_{i}\gamma_{1}^{(i)}\coth{\frac{\hbar\omega_{ge}}{2k_{B}T^{(i)}}},$ (3)
where $n=\left[\exp{\left(\hbar\omega_{ge}/k_{B}T\right)}-1\right]^{-1}$ is the Bose-Einstein distribution.
The population ratio for the qubit then writes
$\frac{p_{e}}{p_{g}}=\frac{\Sigma_{i}\,\Gamma_{\uparrow}^{(i)}}{\Sigma_{i}\,\Gamma_{\downarrow}^{(i)}}=\frac{\Sigma_{i}\,\gamma_{1}^{(i)}n\left(T^{(i)},\omega_{ge}\right)}{\Sigma_{i}\,\gamma_{1}^{(i)}\left[n\left(T^{(i)},\omega_{ge}\right)+1\right]}$ (4)
@@ III Population measurements
In this work we measured population distribution of several transmon qubits to obtain temperature using the technique described in [29], which is based on applying different sequences of $\pi$-pulses and then performing a dispersive readout of the qubit state. Here we do not assume that the transmon is a two-level qubit, but include the higher levels, especially the third one |$\ket{f}$ in the analysis and protocols. We presume, that the population of a qubit thermalized at temperature $T$ follows the Maxwell-Boltzmann distribution $p_{i}=\exp{\left(-E_{i}/k_{B}T\right)}/Z$, with the partition function $Z=\sum_{i=1}^{3}\exp{\left(-E_{i}/k_{B}T\right)}$ and the $i$-th level energy $E_{i}$. A transmon qubit typically has a weak anharmonicity $\alpha$ $(|\alpha|\ll\hbar\omega_{ge})$ [40], so that the energy separations $\hbar\omega_{ge}$ and $\hbar\omega_{ef}$ are related by $\omega_{ef}=\omega_{ge}+\alpha/\hbar\cong\omega_{ge}$. For example, if a transmon with $\omega_{ge}/2\pi=6.65\,$GHz and $\alpha/h=-230\,$MHz is thermalized at temperature 300300300\,300mK we can estimate the population ratios $p_{e}/p_{g}\approx 0.35$ and $p_{f}/p_{g}\approx 0.12$. Thus, it is important to take into account the finite population of the $f$-level. Limiting to the three lowest energy states of a transmon, the density matrix of a thermalized transmon reads
@@
Let us denote the readout response of a projective measurement of the qubit being in state |$\ket{i}$ as $\varphi_{i}$, which we will call “pure state response” further in the text. Then the result $\varphi$ of the qubit state measurement is a linear combination of the form $\varphi=p_{g}\varphi_{g}+p_{e}\varphi_{e}+p_{f}\varphi_{f}$. The $\pi_{ge}$($\pi_{ef}$)-pulses effectively swap pairs of the population probabilities $p_{g}$ and $p_{e}$ ($p_{e}$ and $p_{f}$), which results in getting different readout outcomes, which we denote $x_{i}$ and $y_{j}$, $i,j=0,1,2$. An example of such a measurement is shown in Fig. 1 a, c. We can then write a system of linear equations as
@@
We can find functional dependencies between $p_{i},i\in\{g,e,f\}$ and $x_{j},y_{k}$ with $j,k=0,1,2$ which are valid for arbitrary pure state responses $\varphi_{i}$. These quantities are temperature dependent given by $A=\frac{p_{g}-p_{e}}{p_{g}-p_{f}}$, $B=\frac{p_{e}-p_{f}}{p_{g}-p_{e}}$ and $C=\frac{p_{e}-p_{f}}{p_{g}-p_{f}}$, which can also be expressed through the readout results $x_{j},\ y_{k}$ (see Table 2). To understand the temperature behaviour of functions $A,\ B$ and $C$, we can take the low temperature limit $\hbar\omega_{ge}\ll k_{B}T$, also assuming $p_{f}\rightarrow 0$. Then we can rewrite $A\approx 1-p_{e}/p_{g}$, $B\approx C\approx p_{e}/p_{g}$ or
$A\approx 1-$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)},$ (7)
$B\approx C\approx$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}.$ (8)
IV Sample description
We used transmon qubits on three different chips I, II and III (which is reflected in the name, e.g. R2-I, R4-I, R3-II, Q2-III) made in different fabrication cycles. All transmon qubits had a fixed transition frequency $\omega_{ge}/2\pi\approx 6.5\,$GHz or flux-tunable spectrum ($\omega_{ge}/2\pi=5.5-6.8\,$GHz) and charging energy $E_{c}/h=210-230\,$MHz (see Fig. 1) and they were weakly coupled ($g/2\pi=35-45$ MHz) to a coplanar waveguide (CPWG) $\lambda/4$ readout resonator with the internal quality factor $Q_{\mathrm{int}}\sim 10^{5}$, external quality factor $Q_{\mathrm{ext}}\approx 8\cdot 10^{4}$ and fundamental frequency $\omega_{r}/2\pi=4.5-6.0\,$GHz for a dispersive state readout ($\left|\Delta_{r}\right|=\left|\omega_{ge}-\omega_{r}\right|\gg g$). Experimentally found parameters of the devices are shown in Table 1. The fabrication details are given in Appendix D.
@@ IV Sample description
V Results and discussion

### v3

@@ Abstract
We report temperature measurements using a transmon qubit by detecting the population of its first three energy levels, after applying a sequence of $\pi$-pulses and performing projective dispersive readout. We measure the effective temperature of the qubit and characterize its relaxation and coherence times $\tau_{1,2}$ for three devices in the temperature range of $20-300$ mK. We analyze the process of qubit thermalization to its effective environment consisting of multiple heat baths and support it with experimental data. Signal-to-noise (SNR) ratio of the temperature measurement depends strongly on $\tau_{1}$, which drops at higher temperatures due to quasiparticle excitations, adversely affecting the measurements and setting an upper bound of the dynamic temperature range of the thermometer. The measurement relies on coherent dynamics of the qubit during the $\pi$-pulses. The effective qubit temperature follows closely that of the cryostat in the range of $100-250$ mK. We present a numerical model of the qubit population distribution and compare it favorably with the experimental results. Finally, we compare our technique with previous works on qubit thermometry and discuss its application prospects.
@@ I Introduction
In recent years researchers demonstrated experiments on thermometry of various quantum systems, such as cold gases [1, 2, 3], ion crystals [4], spin ensembles in solutions [5, 6], NV-centers [7, 8, 9, 10, 11], semiconductor quantum dots [12, 13, 14, 15] and superconducting qubits [16, 17, 18, 19, 20, 21]. Quantum thermometry of coherent systems attracts special attention due to prospects of monitoring temperature with minimal invasion, properties of quantum systems with high accuracy at extremely low temperatures. It also provides opportunities for studying open quantum systems and quantum thermodynamics [22, 23, 24, 25, 26, 27, 28, 29, 30, 2, 31, 32, 33, 34, 35, 36, 37, 3, 38, 21].
Superconducting qubits are a well-established platform for realization of quantum electrodynamics (QED) [39] which is widely used for building multi-qubit systems [40, 41, 42, 43]. However, measurements of the population distribution of such qubits are mainly aimed at device characterization and benchmarking of control pulses by analyzing the residual population. This has been achieved using measurements of Rabi oscillations [16, 17], applying different sequences of $\pi$-pulses [20], correlation measurements [18] or single-shot measurements (see for example Ref. [44]). Another approach was demonstrated on the basis of temperature dependent S-parameter of a microwave waveguide directly coupled to a qubit [19] or a two qubit system [21].
@@
Qubits are sensitive to various sources of dissipation and decoherence, and the diverse nature of their environment (see, for example, [45, 42, 46, 47, 48, 49, 50, 51, 52]). Therefore, in practice, identifying which specific thermal bath leads to thermalization of the qubit, can be quite challenging; i.e., it may be difficult to identify what temperature the qubit is actually measuring. Thus, prospects of utilizing superconducting qubits as thermometers are of dual research interest: they can facilitate studies of open quantum systems and interaction between the qubits and their environment, which is vital in the context of building large multiqubit systems such as quantum computers and simulators. On the other hand, qubits can be used for realization of various quantum thermometry protocols.
@@
In this paper we present a comprehensive experimental study of employing a superconducting qubit for quantum thermometry at sub-kelvin temperatures. The measurement technique used in our work relies on applying sequences of $\pi$-pulses for manipulating the qubit population distribution, requiring sufficient coherence of the qubit, and then performing consecutive qubit state readout, which was described in Ref. [20]. We observe and discuss the process of qubit thermalization to its effective environment consisting of different thermal baths, as well as temperature dependencies of qubit relaxation and coherence times. These timescales determine the working range of the thermometer in terms of temperature and introduce constraints on the measurement pulse durations. While showing a solid thermalization of the qubit to the cryostat stage, the technique allows us to measure the population distribution at temperatures up to $200-220$ mK, where the second excited state of the transmon qubit is already significantly populated. This exceeds the applicability range of methods based on consideration of only the lowest transmon levels. At the same time, the technique is less dependent on the signal-to-noise ratio (SNR) of the measurement setup. We compare the experimental results with a numerical model of qubit population measurements. Within the model we investigate the influence of the measured parameters on the population and effective qubit temperature obtained as the result. Finally, we address the factors limiting the dynamical range of a superconducting qubit thermometer and perspective applications. The paper is organized as follows: in Section II we provide a description of a realistic model of a qubit being in thermal equilibrium with several heat baths. Section III explains the principle of the population distribution measurements. Section IV provides a brief description of the samples used in the experiment, measurement sequence, and discussion on the main experimental results. Finally, limits of the thermometer, ways of improvement and possible application, as well as comparison with other experimental works are discussed in Section V. The paper is followed by appendices providing additional experimental data, details of the numerical simulations, experimental setup and sample fabrication.
@@ II Qubit population and effective temperature
Let us consider a qubit in thermal equilibrium with a heat bath having temperature $T$. In this case the qubit is in a thermal state, which means that its density matrix is diagonal and the probabilities $p_{g}$ and $p_{e}$, respectively, in the ground |$\ket{g}$ or excited state |$\ket{e}$ (qubit populations) are Boltzmann distributed as
@@
In a more realistic scenario where a qubit is in a steady state with respect to several uncorrelated heat baths (with temperatures $T^{(i)}$, where $i$ enumerates baths), the state of the qubit is still a thermal-like state. Excitation $\left(\Gamma_{\uparrow}^{(i)}\right)$ and relaxation rates $\left(\Gamma_{\downarrow}^{(i)}\right)$ caused by each bath follow the detailed balance principle
$\frac{\Gamma_{\uparrow}^{(i)}}{\Gamma_{\downarrow}^{(i)}}=\exp{\left(-\frac{\hbar\omega_{ge}}{k_{B}T^{(i)}}\right)}.$ (2)
However, the temperature characterizing such qubit state is not necessarily the temperature $T^{(i)}$ of any of the thermal baths. Then Eq. (1) could be interpreted as giving an effective qubit temperature $T_{\mathrm{eff}}$. In this sense, Eq. (1) can be rewritten as a definition of $T_{\mathrm{eff}}$ as
@@
which coincides with the true temperature in equilibrium. For the total qubit excitation and relaxation rates $\Gamma_{\uparrow}=\Sigma_{i}\Gamma_{\uparrow}^{(i)}$ and $\Gamma_{\downarrow}=\Sigma_{i}\Gamma_{\downarrow}^{(i)}$ the detail balance principle is also valid:
$\frac{\Gamma_{\uparrow}}{\Gamma_{\downarrow}}=\frac{\sum_{i}\Gamma_{\uparrow}^{(i)}}{\sum_{i}\Gamma_{\downarrow}^{(i)}}=\exp{\left(-\frac{\hbar\omega_{ge}}{k_{B}T_{\mathrm{eff}}}\right)}.$ (4)
Then we can introduce a photon number
$n_{\mathrm{eff}}=\frac{1}{\exp{\left(\hbar\omega_{ge}/k_{B}T_{\mathrm{eff}}\right)}-1}=\frac{\sum_{i}\Gamma_{\uparrow}^{(i)}}{\sum_{i}\Gamma_{\downarrow}^{(i)}-\Gamma_{\uparrow}^{(i)}}.$ (5)
If the baths have ohmic spectrum, the rates write
$\Gamma_{\uparrow}^{(i)}=\gamma^{(i)}n_{i}\left(\omega_{ge},T^{(i)}\right),$ (6)
$\Gamma_{\downarrow}^{(i)}=\gamma^{(i)}\left[n_{i}(\omega_{ge},T^{(i)})+1\right],$ (7)
where $\gamma^{(i)}$ are constant coupling rates, which could be obtained from the Fermi golden rule, and $n_{i}\left(\omega_{ge},T^{(i)}\right)=\left[\exp{\left(\hbar\omega_{ge}/k_{B}T^{(i)}\right)}-1\right]^{-1}$ is the Bose-Einstein distribution. The qubit energy relaxation rate then writes
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma^{(i)}$ $\left(2n_{i}+1\right)$
$=\gamma_{1}^{0}\left(2n_{\mathrm{eff}}+1\right)$ $=\gamma_{1}^{0}\coth{\frac{\hbar\omega_{ge}}{2k_{B}T_{\mathrm{eff}}}},$ (8)
where $\gamma_{1}^{0}=\sum_{i}\gamma^{(i)}$ is the total relaxation rate at zero temperature. Finally, we can rewrite $n_{\mathrm{eff}}$ from Eq. (5) as
$n_{\mathrm{eff}}=\frac{1}{\gamma_{1}^{0}}\sum_{i}\gamma^{(i)}n_{i},$ (9)
which can be considered as the average photon number over different baths, weighted by their coupling coefficients.
@@ III Population measurements
In this work we measured the population distribution of several transmon qubits to obtain their effective temperature using the technique described in Ref. [20], which is based on applying different sequences of $\pi$-pulses and then performing a dispersive readout of the qubit state. We consider a qutrit model, which accounts for population of three lowest energy levels in the analysis and protocols. We presume, that the population of a qubit thermalized at temperature $T$ follows the Maxwell-Boltzmann distribution $p_{i}=\exp{\left(-E_{i}/k_{B}T\right)}/Z$, with the partition function $Z=\sum_{i=1}^{3}\exp{\left(-E_{i}/k_{B}T\right)}$ and the $i$-th level energy $E_{i}$. A transmon qubit typically has a weak anharmonicity $\alpha$ $(|\alpha|\ll\hbar\omega_{ge})$ [53], so that the energy separations $\hbar\omega_{ge}$ and $\hbar\omega_{ef}$ are related by $\omega_{ef}=\omega_{ge}+\alpha/\hbar\cong\omega_{ge}$. For example, if a transmon with $\omega_{ge}/2\pi=6.65\,$GHz and $\alpha/h=-230\,$MHz is thermalized at temperature 300300300\,300mK we can estimate the population ratios $p_{e}/p_{g}\approx 0.35$ and $p_{f}/p_{g}\approx 0.12$. Thus, it is important to take into account the finite population of the $f$-level. Limited to the three lowest energy states of a transmon, the density matrix of a thermalized transmon reads
@@
Let us denote the readout response of a projective measurement of the qubit being in state |$\ket{i}$ as $\varphi_{i}$, which we will call “pure state response” further in the text. Then the result $\varphi$ of the thermalized qubit state measurement is a linear combination of the form $\varphi_{\mathrm{therm}}=p_{g}\varphi_{g}+p_{e}\varphi_{e}+p_{f}\varphi_{f}$. From the experimental point of view, $\varphi$ here is a measured complex valued voltage. To measure it we send a readout pulse with a rectangular envelope at a frequency close to the readout resonator frequency $\omega_{r}$, which is transmitted through the sample waveguide, then it is amplified and downconverted. Finally, it is measured by a fast (2 GS/s) analog-to-digital converter (ADC) and integrated over the pulse duration time, which is typically $0.1-2.0\,\mu$s.
However, a single measurement of $\varphi_{\mathrm{therm}}$ is not enough to extract the populations $p_{i}$ as the pure state responses $\varphi_{i}$ are unknown. To solve this problem, before the qubit state readout is performed, $\pi_{ge}$- and $\pi_{ef}$-pulses can be applied at frequencies $\omega_{ge}$ and $\omega_{ef}$, respectively, that leads to swapping of the population between |$\ket{g}$ and |$\ket{e}$ states or |$\ket{e}$ and |$\ket{f}$ states, correspondingly. This results in different readout signals, which we denote $x_{j}$ and $y_{k}$, $j,k=0,1,2$. An example of such a measurement is shown in Fig. 1 a, c. For different $\pi$-pulse sequences we can then write a system of linear equations as
@@
Ratio Calculation method Result Equation
1 2 3
@@
We can find functional dependencies between $p_{i},i\in\{g,e,f\}$ and $x_{j},y_{k}$ which are valid for arbitrary pure state responses $\varphi_{i}$. These quantities are temperature dependent given by $A=\frac{p_{g}-p_{e}}{p_{g}-p_{f}}$, $B=\frac{p_{e}-p_{f}}{p_{g}-p_{e}}$ and $C=\frac{p_{e}-p_{f}}{p_{g}-p_{f}}$, which can also be expressed through the readout results $x_{j},\ y_{k}$ (see Table 1). To understand the temperature behavior of functions $A,\ B$ and $C$ better, we can take the low temperature limit $\hbar\omega_{ge}\ll k_{B}T$, where $p_{f}\rightarrow 0$. Then we can rewrite $A\approx 1-p_{e}/p_{g}$, $B\approx C\approx p_{e}/p_{g}$ or
$A\approx 1-$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)},$ (12)
$B\approx C\approx$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}.$ (13)
We extract the effective qubit temperature $T_{\mathrm{eff}}$ by numerically solving equations for functions $A,B$ and $C$, shown in Tab. 1.
Population ratios $A,\ B$ and $C$ can be extracted from the measurement data $x_{j},\ y_{k}$ by three different methods each, which we denote as $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$ (see Tab. 1). Qualitatively speaking, the difference between the methods is the order of the $\pi$-pulses in the control sequence. Because of the qubit relaxation and relatively low population of the $f$-level, the resulting signals can have different signal-to-noise (SNR) ratio, giving us some guidance on how to choose the best one (see discussion in Section V and Appendix C).
IV Experimental details
We used transmon qubits on three different chips I, II and III (which is reflected in their names: R2-I, R4-I, R3-II, Q2-III) made in different fabrication cycles. Each transmon qubits had either fixed transition frequency $\omega_{ge}/2\pi\approx 6.5\,$GHz or flux-tunable spectrum ($\omega_{ge}/2\pi=5.5-6.8\,$GHz) and charging energy $E_{c}/h=210-230\,$MHz (see Fig. 1) and they were weakly coupled ($g/2\pi=35-45$ MHz) to a coplanar waveguide (CPWG) $\lambda/4$ readout resonator with the internal quality factor $Q_{\mathrm{int}}\sim 10^{5}$, external quality factor $Q_{\mathrm{ext}}\approx 8\cdot 10^{4}$ and fundamental frequency $\omega_{r}/2\pi=4.5-6.0\,$GHz for a dispersive state readout ($\left|\Delta_{r}\right|=\left|\omega_{ge}-\omega_{r}\right|\gg g$). Experimentally found parameters of the devices are shown in Table 2. The fabrication details are given in Appendix D.